OpenStudy (anonymous):

During the day, shadows can be used to find the height of items that are too tall to measure using standard measuring tools. Explain how to use shadows to indirectly measure the height of a tall object. Be specific in your explanation including all the steps that would need to be taken to calculate the height.

OpenStudy (anonymous):

hey did you read ur messages

OpenStudy (anonymous):

ear King George III of Great Britain, I am writing to you because ______________________________. I write only after long thought about the many issues between us. Some issues and events of the last few years include _______ _______________________. You see, in the American colonies, we believe that the purpose of government is _____________ _________________. Government should ______________________________. It should follow the principles of John Locke, such as ______________________________ and __________ ____________________. These ideas mean that ______________________________. The role of government in natural law is to ______________________________. Unfortunately, you have not lived up to these expectations as a ruler. One example of how you have disappointed us is ______________________________. This affects our rights because______________________________. Another example is ___________ ___________________. This affects our rights because______________________________. A third example is ______________________________. This affects our rights because______________________________. Please remember that we did try to reason with you first. Remember when we _____________ _________________, and then you ______________________________? Or, there was the time we ______________________________. Because you have failed to address our concerns, the American colonists are declaring complete independence from you and from your country, Great Britain. We are no longer ______________________________ and from now on, we will

OpenStudy (anonymous):

heeeeeeeeeeeeeeeeeeeeelllllllllllllllp please

OpenStudy (anonymous):

You can measure the height of a tall object indirectly by using the length of its shadow. At the same time of day, measure the tall objects shadow and also measure the shadow of an object of known height. The shadows are the bases of similar triangles; the heights are proportional. Example: the shadow of an object 10m high is 5m, at the same time the shadow of an object of unknown height (x) is 20m; 10/5 = x/20, 5x = 200, x = 40m.

OpenStudy (anonymous):

idk ig u would use geometrical equations

OpenStudy (anonymous):

since it is shapes and measurements